A horizontal pipe contains water at a pressure of 140 kPa flowing with a speed of 2.0 m/s. The pipe narrows to one-half its oPart A What is the speed of the water? Express your answer using two significant figures. ΟΙ ΑΣΦ ? V= m/sPart B What is the pressure of the water? Express your answer using two significant figures. ΥΠΙ ΑΣφ ? P = kPa
Answer #1

A.

V = V1/(½)² = 2/(1/4) = 8 m/s

B.


P = P1 - ½ρ(V² - V1²) = 140E3 - .5*1000*(8² - 2²) = 110 KPa